OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = 2x^2 + 4x – 6? x–intercepts: (1, 0) and (–3, 0) y–intercept: (0, –3) x–intercepts: (1, 0) and (–3, 0) y–intercept: (0, –6) x–intercepts: (–1, 0) and (3, 0) y–intercept: (0, –3) x–intercepts: (–1, 0) and (3, 0) y–intercept: (0, –6)

OpenStudy (anonymous):

x–intercepts: (1, 0) and (–3, 0) y–intercept: (0, –6)

OpenStudy (anonymous):

the x-intercepts is where y= 0 plug y = 0 to the equation y = 2x^2 + 4x – 6 0 = 2x^2 + 4x – 6 divide all by 3 -> 0 = x^2 + 2x - 3 0 = (x+3)(x-1) x = -3 or x = 1

OpenStudy (anonymous):

the y-intercepts is where x =0 plug x = 0 to the equation y = 2x^2 + 4x – 6 y = -6

OpenStudy (anonymous):

therefore the answer is x-intercepts : (1,0) and (-3,0) y intercept : (0,-6)